# Training Bilingual LMs with Data Constraints in the Targeted Language

### v1

@@ 4.2 Clustered Dataset Importance Sampling
Prior work shows that LLMs reason in English and that information may be stored in a language agnostic space Wendler et al. (2024). However, they do not control how the information is seen during training, and while much of the information may be seen only in English (the predominant language), it may also be seen in other languages as the pretraining datasets are not made publicly available. Further, we note that FineWeb-EDU, in addition to high quality filtering, also filters for educational quality content Penedo et al. (2024).
@@
To explicitly test whether information is shared between auxiliary and target languages, we upsample topics in English and evaluate on the target language comparing with having uniform data. Specifically, given access to some small dataset representative of target knowledge, we group the data into target clusters and estimate importance sampling weights over the clusters following Grangier et al. (2024b). To train the clustering model, we take a small subset of the training set, produce embeddings from a smaller sentenceBERT model, then cluster the data according to the embeddings. For our importance sampling experiments, we upweight a subset of roughly 300B tokens from the English dataset. Given a small target set (typically on the order of 1000-10000 samples), we assign each sample to a cluster and upweight the original training set based on the cluster assignment proportions. In practice, we do not optimize for the cluster parameters jointly with the model weights, and instead precompute them based on the pretraining and target task data.
For clustering hyperparameters, we use a lightweight SentenceTransformers multilingual model [^3] The particular model is called paraphrase-multilingual-MiniLM-L12-v2 model and is obtained from https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2. for extracting features Reimers and Gurevych (2019), and a balanced $K$-means algorithm to cluster the embeddings into 64 clusters. To learn importance sampling weights, we use corresponding training sets, and available corpora that would be used for retrieval-based methods on the same tasks.
All upsampling is done in English to facilitate having specialized information in the auxiliary language, but not in the target language to measure the impact of information in a different language. For all experiments, we use the mC4 English dataset. Comparisons are made between having access to more German data of uniform quality, upsampling based on the ARC training set (general science knowledge), and upsampling based on the HellaSwag training set (general knowledge and QA style).
@@
Results in Figure 4 show that for the medium model experiments, data selection provides 1.5% improvement over base mC4 auxiliary data on English language evaluations, but no improvement in the target language (German). For the XL models, we see 4% improvement in English evaluations, and 2% improvement in target language (German). These results highlight that larger models can take advantage of information in the auxiliary language, and the performance improvements are higher for the information upsampling than for model-based filtering without removing the data, allowing for training on new data at greater token counts.
(a) Medium
(b) XL
Figure 5: Zero-shot accuracy of medium and XL models trained upsampling based on synthetic datasets. Results are averaged over six eval datasets.
Our findings indicate that while filtering for high quality data as in the OH filter has limited improvement for the target language, but further improvements come from data selection over important topics.
@@ 4.3 Upsampling with Synthetic Examples
While performance improvements are achieved by data selection based on target downstream evaluations, having such data available at pretraining can be restrictive. For this reason, it can be desirable to be able to generate the necessary data for upsampling. While prior work has examined the use of LLM-generated data for pretraining Maini et al. (2024) and finetuning Li et al. (2023a); Yuan et al. (2024) language models, to our knowledge there is no prior work that investigates data selection based on synthetic examples.
@@
We generate a small set of synthetic examples following the approach in Maini et al. (2024). The synthetic data is created by prompting an off-the-shelf instruction finetuned language model to generate sets of questions relating to the topic of interest using the prompts in Section D. Generating synthetic data using an off-the-shelf language model can be both computationally expensive and challenging, however for the purpose of computing sampling weights, a small number of questions is sufficient.
For our experiments, we generate data using a frozen Mistral-7B instruction tuned model [^4] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 Jiang et al. (2023). In total around 2000 science question answer pairs (referred to as SciQ) are created with minimal filtering outside of confirming that both a question and answer are specified, and the model does not generate any additional details or explanations.
We additionally generate general instruction data. This data is aimed at broad QA style and general fact information which is helpful for downstream tasks. This data is generated in two stages by first having the frozen Mistral-7B model generate a set of questions, then answering the questions.
Using the sets of questions, we identify the cluster sample weights and upsample data accordingly as in Section 4.2. The approach otherwise follows the same setup as in Section 4.2 except for the synthetic QA pairs.
@@
Results shown in Figure 5 comparing synthetic data upsampling with upsampling from the downstream task demonstrates that synthetic data can be sufficient for incorporating information into the auxiliary English dataset. For the medium models on both English and German evaluations, models trained by upsampling synthetic data and real ARC attain the same performance, and for the XL models, upsampling synthetic data is within 1% for the English downstream tasks, and within 0.5% for the translated German evaluations.
@@ 4.4 Translation Systems
(a) Medium
(b) XL
Figure 6: Zero-shot accuracy of medium and XL models trained upsampling based on translating mC4. Results are averaged over six eval datasets.
@@
Training directly on auxiliary language data can lead to improvements. However, an alternative strategy is to translate the auxiliary data into the target language, assuming a machine translation system is available between the two languages. This approach offers the benefit of training the model exclusively in one language, and, if the translation system is of high quality, it allows for training on high-quality data in the target language at the expense of translating the corpus. It is, therefore, crucial to also assess the level the translation system must possess to effectively translate data for pretraining.
@@
For our experiments, we use light-weight translation systems of roughly 100-200M parameters. We consider three models with BLEU scores 16.0, 26.5, and 31.6 on the WMT-17 EN-DE benchmark task. We denote these models as v1, v2, and v3 corresponding to increasing BLEU score. All models are trained on translated versions of the mC4 english corpus. No other English data is included in the dataset, but we keep the 250M tokens of real German data as in prior experiments.
@@
We summarize our results in Figure 6. For all translation models, we find little difference from the quality of the translation. Both stronger translation models achieve similar performance with the worst model performing comparably to training on real German data. For the medium models, we see that the performance on the translated German evaluations yields similar performance to training on real German data. For the XL models, the translated data improves by around 1% over new German data for the v2 and v3 translation models. We hypothesize a few reasons this may be possible, but leave investigation of each of these to future work: (1) English CC data is higher quality than German CC data as there may be more data from more diverse sources. (2) Translated German data has a different distribution from real German data and this better matches the translated test evaluations. (3) Translated data from a small translation system might simplify language, making it easier for models to learn, following Eldan and Li (2023). (4) Portions of the dataset could not be translated by the systems and are removed. These portions might be noisy, and some unintended filtering may lead to slightly higher performance.
5 Do Findings Hold Across Multiple Languages?
Sections 3-4 focused on German as the target language. We now investigate to what extent our findings for German hold across different target languages. In Section 5.1 we study additional languages individually, and in Section 5.2 we investigate a multilingual setting combining a subset of four languages.
5.1 Experiments Across Multiple Languages
@@
We add seven languages, across four language families: French, Italian, Portuguese, Spanish (Indo-European, same as German), Chinese (Sino-Tibetan), Japanese (Japonic), and Korean (Koreanic) Lewis et al. (2015).
@@
For these experiments, we train models using the same experimental setup as in Section 3. For all additional target language models, we train with approximately 250M tokens from the mC4 corpus in the respective language. For training monolingual models in the target language, we note that the Chinese and Korean mC4 corpora contain fewer than 100B tokens and thus the data is repeated for multiple epochs in the base language experiments.
@@
We report results for the XL models trained for 100B tokens in Figure 10-13 in the appendix. Results are similar for the medium models. Our results indicate that our findings on English-German training do not extend across all families of languages. In particular, we see improvements from FineWeb-EDU only on the Indo-European languages.
We investigate the performance differences further by evaluating the perplexity of models trained on combinations of target language data from mC4 and FineWeb-EDU. We selected four languages (German, French, Chinese, and Japanese) to examine the perplexity across two languages with improved performance from auxiliary data and two languages without. We compare data from the FineWeb-EDU dataset seen during training with data from the mC4 English portion which has not been seen during training. We translate 10,000 documents into each of the target languages and evaluate the perplexity of documents for both translated and original auxiliary data. We additionally report a metric we call exceedance, which is the fraction of documents in FineWeb-EDU with lower perplexity than the average perplexity of documents in mC4 English. A higher value shows that the data the model is trained on appears more in-distribution, and similar values across original and translated documents would reflect transfer between the languages. We observe that the exceedance is similar for French and German between translated and original documents (Table 3, Appendix F.2). However for Japanese and Chinese [^5] For Chinese translations, there may be some artifacts that induce artificially high perplexity in the original Chinese mC4 data and further decrease exceedance. In particular, we find several symbols and characters that are not removed in text extraction. Nonetheless we believe the patterns should be consistent from Japanese. translations, the exceedance is lower, and translated FineWeb-EDU data has the same or higher perplexity compared with mC4 English data.
5.2 Multilingual Experiments
Motivation.
Methodology.
We conduct experiments combining the German, French, Chinese, and Japanese language data towards multilingual training. For these experiments, we train a 1B model with the auxiliary dataset being FineWeb-EDU, and a mix of data from the four languages totaling 5% or 20% of the training. We again use the four language subset as we keep the data ratios the same per language, and did not want to increase the amount of data in target languages beyond the typical multilingual ratios in large open-source models Xu et al. (2024).
Findings.
EN DE FR JA ZH
5% 61.07 46.02 46.64 44.00 45.92
20% 59.25 46.44 46.61 43.23 44.36
Bi 47.16 47.52 42.73 44.38
Table 1: Evaluation of XL models in multilingual setting on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in translated languages. Rows are the average accuracy for the respective language, with 5% or 20% of the training coming from a mix of the four languages. ‘Bi’ refers to the bilingual models.
6 Scaling Limitations for Low Resource Languages
We consider the limitations of data scaling for training bilingual language models for target languages with limited data. In Section 6.1, we investigate the extent to which training on higher quality English data improves the amount of data needed in the target language to reach similar performance. In Section 6.2 we investigate how feasible it is to train larger bilingual models with limited target data.
6.1 Auxiliary Data Scaling
Figure 7: Average accuracy over zero-shot benchmark tasks in translated German with increasing number of tokens in both target and auxiliary languages. Models are XL size and trained for 100B tokens.
(a) 1B Model Perplexity
(b) 3B Model Perplexity
Motivation.
We measure the amount of data needed in the target language to match training on the high quality auxiliary data. Our goal is to quantify the advantage of training models with additional target data beyond the 250M tokens used in prior experiments. For this experiment we train models at different dataset sizes ranging from 0.1B to 100B tokens.
Methodology.
We investigate data scaling over both the mC4 dataset and FineWeb-EDU to measure the increase in German data needed to match higher quality auxiliary datasets.
Findings.
@@ 6.1 Auxiliary Data Scaling
Our results are summarized in Figure 7. Models trained on the base mC4 English data achieve similar performance to models trained with only 5B tokens of the base mC4 German data. In contrast, training on a high quality data such as FineWeb-EDU increases the amount of data needed to around between 10-15B tokens or equivalently 2x the amount of data in German. Note that the curves all plateau quickly at around 10B tokens. This corresponds to around 10 repetitions of the data, matching the results in Muennighoff et al. (2024).
@@
There are two important data scaling considerations from Figure 7 that justify training with auxiliary language data. First, for languages that are data constrained, it may be infeasible to collect twice as much data. Second, models trained on FineWeb-EDU attain similar performance at a rate of 5x the number of tokens (roughly 50B tokens on FineWebEDU matches the performance of 10B tokens of mC4 German data). An important avenue of research is to investigate data scaling at larger quantities of tokens. In particular, the FineWeb-EDU corpus totals 5.4T tokens and would require access to 1T tokens of German data, which is 3x the amount of data in mC4. As a result, while the data scaling shows large improvements from little German data, the large amount of readily available English data can make training on auxiliary data practical.
(a) 250M Tokens
(b) 1B Tokens
Figure 9: Zero-shot accuracy for 1B and 3B parameter models trained with either 250M tokens of data in the target language or 1B tokens. Both models are trained on the target language data for 5% of the training steps.
@@ 6.2 Model Size Data Scaling
In this section, we investigate to what extent results scale for training larger language models when data remains constrained in the target language.
Motivation.
In prior experiments, we fixed the amount of data from the target language to approximately 250M tokens and train for 100K steps with 5% of the training being in the target language. However, increasing model size necessitates increasing the number of tokens seen during training according to Chinchilla scaling laws Kaplan et al. (2020). This is challenging for low resource languages, as the number of repetitions increases with increased amount of training, which can lead to overfitting and saturate model performance.
Methodology.
We first examine the amount of overfitting larger models have on the same amount of data and ratio of training time. We train a roughly 3B (non-embedding parameter) model consisting of 32 layers, 32 attention heads, and a hidden dimension size of 2560. The model has a maximum sequence length of 2048 and is trained for 150K steps to match the same scaling ratio as for the 300M and 1B models.
Findings.
The perplexity of training and validation data is shown in Figure 8 for the 1B (left) and 3B models (right). For 1B models, we see little to no overfitting at 5% ratio of training steps and the model achieves slightly lower perplexity than a model trained with a lower ratio of 1.5% target language data. In contrast, for the 3B model, there is clear overfitting from as early as 25% of the training with 5% ratio of training steps. This indicates that the number of repetitions is too high for the 3B model and performance may degrade.
Reducing the number of repetitions by reducing the fraction of training steps to 1.5% reduces overfitting as seen in Figure 8(b), however perplexity remains similar between the 1B and 3B models when limited to 250M tokens. Increasing the number of available tokens results in decreasing perplexity on validation data and no overfitting for the 3B model, which is unnecessary for the 1B models. An overview of comparison with different data ratios for zero-shot accuracy is also provided in Appendix E.
Based on Figure 8, we evaluate both 1B and 3B models with data ratios that mitigate overfitting: 5% for 1B and 1.25% for 3B with 250M tokens in the target language. Results are shown in Figure 9. First, Figure 9(a) shows that the 3B model performs similarly to the 1B model with the same 250M amount of target language data matching the similar perplexity values between the two models. However increasing to 1B tokens results in similar improvements in both English and German when increasing model size from 1B to 3B as shown in Figure 9(b).
Summarizing, there is a limit to the size of models that can be trained with limited target data. Model performance for 250M tokens saturates at 1B parameter models (more than the available data in mC4 for $\sim 20$ languages). Increasing to a 3B parameter model necessitates training on 1B tokens of data (more than the available data in mC4 for $\sim 40$ languages).
@@ 7 Conclusion
This work studies how an auxiliary language for which an abundance of training data is available can boost pretraining for a target language for which only limited data is available. We find that adding auxiliary high quality data obtained by data filtering can improve performance in a target language. Moreover, our results indicate that most of these gains can be attributed to having relevant information in the auxiliary language data, which may not be present in the target language. However, we find that results are inconsistent across different target languages. We hypothesize that for languages that are further from English, better English datasets are not as helpful as information is not shared between them. Finally, we find limitations to scaling models for languages that are data constrained. This work takes a step towards pretraining language models in languages with limited data, and can inspire more research into bilingual or multilingual learning under dataset constraints.

### v2

@@ 4.2 Clustered Dataset Importance Sampling
Prior work shows that LLMs reason in English and that information may be stored in a language agnostic space Wendler et al. (2024). However, it is unclear whether the relevant information is only seen in English (the predominant language), or also in other languages. Our experiments on FineWebEDU indicate the presence of information sharing, but FineWebEDU filters for both high quality data and for educational quality content Penedo et al. (2024). In this section we experiment with isolating the information sharing.
@@
To explicitly test whether information is shared between auxiliary and target languages, we upsample topics in English and evaluate on the target language. Given access to some small target set, importance sampling weights are computed based on the amount of data from a target evaluation set that is assigned to each cluster following Grangier et al. (2024b).
To train the clustering model, we take a small subset of the training set, produce embeddings from a smaller sentenceBERT model Reimers and Gurevych (2020), and cluster the data according to the embeddings. We upweight a subset of roughly 300B tokens from the English dataset. Given a small target set (typically on the order of 1000-10000 samples), we assign each sample to a cluster and upweight the original training set based on the cluster assignment proportions. In practice, we do not optimize for the cluster parameters jointly with the model weights, and instead precompute them based on the pretraining and target task data.
For clustering hyperparameters, we use a lightweight SentenceTransformers multilingual model [^2] The particular model is called paraphrase-multilingual-MiniLM-L12-v2 model and is obtained from https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2. for extracting features Reimers and Gurevych (2019), and a balanced $K$-means algorithm to cluster the embeddings into 64 clusters.
Upsampling is done in English at the cluster level, to facilitate having specialized information in the auxiliary language that is unavailable in the target language. We compare two settings: (i) upsampling based on HellaSwag (general knowledge and instruction style), and (ii) upsampling based on ARC (general science knowledge). We also add a comparison where we downsample relevant data to the target task, to compare with having only low quality data, which we call “no ARC DE (L)”.
Figure 6: Zero-shot accuracy of models trained with upsampling of English auxiliary data. Results are averaged over six evaluation datasets. For each setting evaluation is done in English and German.
@@
Figure 6 shows the results. We see 4% improvement in English evaluations, and 2% improvement in the target language. This highlights that models can take advantage of information in the auxiliary language, and the performance improvements are higher for the information upsampling than for model-based filtering.
Figure 7: Zero-shot accuracy of models trained with synthetic data upsampling of English auxiliary data. Results are averaged over six evaluation datasets. For each setting evaluation is done in English and German.
Generally, our findings indicate that filtering for high quality data has limited improvement for the target language (0.8% averaged across all languages), and further improvements come from data selection over important topics.
@@ 4.3 Upsampling with Synthetic Examples
While we achieved performance improvements by selecting data based on target downstream evaluations, having such data available at pretraining can be restrictive. For this reason, it can be desirable to be able to generate the necessary data for upsampling. While prior work has examined the use of LLM-generated data for pretraining Maini et al. (2024) and finetuning Li et al. (2023a); Yuan et al. (2024) language models, to our knowledge there is no prior work that investigates data selection based on synthetic examples.
@@
We generate a small set of synthetic examples, following the approach in Maini et al. (2024). The synthetic data is created by prompting an off-the-shelf instruction finetuned language model to generate sets of questions relating to the topic of interest. We use the prompts in Appendix D. Generating synthetic data using an off-the-shelf language model can be both computationally expensive and challenging. However, for the purpose of computing sampling weights, a small number of questions is sufficient.
For our experiments, we generate data using a frozen Mistral-7B instruction tuned model [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 Jiang et al. (2023). We create around 2000 science QA pairshile with filteog that cnfirms only a question and answer are specified. We refer to this setting as “S EN (L)”, S referring to Science.
We also generate general instruction data, aimed at broad QA style data, and general factual information helpful for downstream tasks. This data is generated in two stages: we use the frozen Mistral-7B model to generate a set of questions, and then the model is prompted to answer the questions.
Using the sets of questions, we identify the cluster sample weights and upsample data accordingly as in §4.2. We refer to this setting as “S+I EN (L)”.
@@
We show results in Figure 7. Comparing synthetic data upsampling with upsampling from the downstream tasks (Figure 6) demonstrates that synthetic data can be sufficient for incorporating information into the auxiliary English dataset. Upsampling synthetic data is within 1% for the English downstream tasks, and within 0.5% for the translated German evaluations. Results for transformations in other languages are reported in §H.4.
@@ 4.4 Translation Systems
Training directly on auxiliary language data can lead to improvements. An alternative strategy is to translate the auxiliary data into the target language, assuming a machine translation system is available. This approach offers the benefit of training the model exclusively in one language, and, if the translation system is of high quality, it allows for training on high-quality data in the target language at the expense of translating the corpus. It is, therefore, important to also assess the level the translation system must possess to effectively translate data for pretraining.
@@
For our experiments, we use light-weight translation systems of roughly 100-200M parameters. We consider three models with BLEU scores 16.0, 26.5, and 31.6 on the WMT-17 EN-DE benchmark task. We denote these models as v1, v2, and v3, corresponding to increasing BLEU score. All models are trained on translated versions of the mC4 English corpus. No other English data is included in the dataset, but we keep the 250M tokens of real German data as in prior experiments.
Figure 8: Zero-shot accuracy of models trained with translations of English auxiliary data. Results are averaged over six evaluation datasets. For each setting evaluation is done in English and German.
@@
Figure 8 shows the results. We find little difference from the quality of the translation. Both stronger translation models achieve similar performance. The v1 model performs comparably to training on real German data, and the data translated by the v2 and v3 models obtain around 1% improvement in comparison to real German data.
We hypothesize a few reasons this may be possible, but leave investigation of each of these to future work: (1) English CC data is of higher quality than German CC data, as there may be more data from more diverse sources. (2) Translated German data has a different distribution from real German data, and this better matches the translated test evaluations. (3) Translated data from a small translation system might simplify language, which makes it easier for models to learn, following Eldan and Li (2023). (4) Portions of the dataset could not be translated by the systems and are removed. These portions might be noisy, and some unintended filtering may lead to slightly higher performance.
5 Data and Model Scaling
To study the impact of data and model scaling, we investigate two scaling questions: (1) How does the required quantity of additional auxiliary data relate to the amount of data required if one had enough target language data to train on? (§5.1) (2) How do findings extend to a larger bilingual model? (§5.2) Following our setup in the previous section, we target our investigation to German (target) and English (auxiliary).
5.1 Auxiliary Data Scaling
@@
We measure the amount of data needed in the target language to match training on high quality auxiliary data. Our goal is to quantify the advantage of training models with additional target data, beyond the 250M tokens used before.
@@
We investigate data scaling over mC4 and FineWebEDU. We train models at different dataset sizes, from 0.1B to 100B tokens.
Figure 9: Average accuracy over zero-shot benchmark tasks in translated German, with increasing number of tokens in both target and auxiliary languages. Models are trained for 100B tokens.
@@
Our results are summarized in Figure 9. Models trained on the English mC4 data achieve similar performance to models trained with only 5B tokens of German mC4 data. In contrast, training on a high quality data such as FineWebEDU increases the amount of German data needed to around 10B tokens, 2x the amount of German data. Note that the curves all plateau quickly at around 10B tokens, corresponding to around 10 repetitions of the data, matching Muennighoff et al. (2024).
@@
There are two important data scaling considerations from Figure 9 that justify training with auxiliary language data. First, for languages that are data constrained, it may be infeasible to collect twice as much data. Second, models trained on FineWebEDU attain similar performance at a rate of 5x the number of tokens (roughly 50B tokens of FineWebEDU matches the performance of 10B tokens of mC4 German data). An important avenue is to investigate data scaling at larger quantities of tokens. In particular, the FineWebEDU corpus totals 5.4T tokens and would require access to 1T tokens of German data, which is 3x the amount of data in mC4. As a result, while the data scaling shows large improvements from little German data, the large amount of readily available English data can make training on auxiliary data practical.
@@ 5.2 Model Size Data Scaling
Figure 10: Zero-shot accuracy for 1B and 3B models trained with 250M or 1B tokens in the target language denoted in brackets. Models are trained on the target language data for different ratios of the training steps . See §F for details on data ratios.
Next, we investigate to what extent results scale for training larger language models when data remains constrained in the target language. Given a fixed amount of data from the target language (250M tokens) training a larger model may be impractical, as, according to Chinchilla scaling laws Kaplan et al. (2020), increasing model size necessitates increasing the number of tokens seen during training. This is challenging for data constrained languages, as the number of repetitions increases with increased amount of training. This can lead to overfitting and saturated model performance.
We train a $\sim$2.7B parameter model for 150K steps, matching the same scaling ratio as for the 1.3B model. We train models following the DE (S) + EN (L) setup. Our findings are in Figure 10. There is a limit to the size of models that can be trained with limited target data. Model performance for 250M tokens (more than available in mC4 for $\sim$20 languages) shows limited improvement when scaling up from a 1.3B to a 2.7B parameter model. In contrast, increasing to 1B tokens doubles improvement when increasing model size to a 2.7B parameter model. Training a 2.7B parameter model necessitates training on at least 1B tokens of data to see larger gains in performance (more than available in mC4 for $\sim$40 languages).
@@ 6 Conclusion
This work studies how an auxiliary language for which an abundance of training data is available can boost pretraining for a target language for which only limited data is available. We find that adding auxiliary high quality data obtained by data filtering can improve performance in a target language. Moreover, our results indicate that most of these gains can be attributed to having relevant information in the auxiliary language, which may not be present in the target language. However, we find that results are inconsistent across target languages. We hypothesize that for languages further from English, better English datasets are not as helpful as information is not shared between them. Finally, we find limitations to scaling models for languages that are data constrained. This work takes a step towards pretraining language models in languages with limited data, and can inspire more research into bilingual learning under data constraints.